OpenStudy (anonymous):

An outdoor storage bin contains 15,000 kilograms of road salt. Each day, a quarter of the salt is removed. How much salt is in the bin at the end of the 18th day?

OpenStudy (anonymous):

your just going to have to divide 15,000 into 4 then multiply by three and repeat the same process 17 more times

OpenStudy (anonymous):

Why would I divide by 4? Why would I multiply it by 3? The way I was doing it is I multiplied 15,000 by .25 and got 3,750. So, I subtracted that from 15,000 and got 11,250. Then I multiplied 11,250 by .25 and got 2,812.5. So, I subtracted that from 11,250 and got 8,437.5. I would keep doing that until I reached 18, but that would take way too long. I'm not even sure if I'm doing it correctly.

OpenStudy (anonymous):

It's much easier just to divide and multiply and the answer is more certain the correct answer is 63.4242399

OpenStudy (anonymous):

Okay, thank you.

OpenStudy (anonymous):

Anytime :)

OpenStudy (wolf1728):

blueberry520 Basically, this should be approached like a "half life" problem. The difference here is that for every day, the amount remaining is one quarter (instead of one half). So, the remaining amount equals: 15,000 *(.75^t) where 't' is the time in days. So after 2 days the amount remaining is 15,000*(.75^2)= 15,000*0.5625 which equals 8,437.50 after 2 days After 3 days, the amount remaining 15,000*(.75^3)= 15,000*0.421875 which equals 6,328.125 So, at the end of the eighteenth day the amount remaining equals 15,000*(.75^18)= 15,000*0.0056377101136604 which equals 84.5656517049 kilograms By the way, Shelby's answer is correct for the NINETEENTH day. Mine is correct for the 18 days.